OpenStudy (anonymous):

A function is given. Determine the average rate of change of the function between the given values of the variable. f(x) = x + x4; x = −4, x = 3

OpenStudy (anonymous):

sorry it should be f(x)= x+ x^4

OpenStudy (anonymous):

anyone can help?

OpenStudy (anonymous):

Average rate of change of a function f of an interval [a,b] is given by \[\frac{f(b)-f(a)}{b-a}\] In this case, the interval is (-4,3).

OpenStudy (anonymous):

so should I graph this?

OpenStudy (anonymous):

I tried pluging in and I ended up with 171/7

OpenStudy (anonymous):

thats with b=3 and a =-4 like it should

OpenStudy (anonymous):

The average value is \[\frac{f(3)-f(-4)}{3-(-4)}=\frac{\left(3+3^4\right)-\left(-4+(-4)^4\right)}{3+4}\] As far as I can tell, you don't have to do any graphing.

OpenStudy (anonymous):

ya thats what I did

OpenStudy (anonymous):

I messed up, I did it again and I got 168/7 but thats still wrong

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

The numerator can be kinda tricky: \[(3+3^4)-(-4+(-4)^4)\\ (3+81)-(-4+256)\\ 84-252\\ -168\] And divided by 7, you have -24.

OpenStudy (anonymous):

ya I forgot to put the negative im sorry thanks again but u know what it is, I dont put it out in steps like that like I should so I completely got wrong answers the first couple times because of the sign change from the minus. THe way I normally do it is I do one tyerm first on a seperate part of the paper then the other like under it or something and then bring everything down.